(B)Find equilibrium Price and quantity from the following demand and supply functions.

          I.           P=80-Q(demand)                                     P=20+2Q (supply)



1
Expert's answer
2021-03-26T11:59:08-0400

The market equilibrium occurs when "P_d=P_s":


"80-Q_E=20+2Q_E,""60=3Q_E,""Q_E=20."

Finally, we can find the equilibrium price:


"P_E=80-20=\$60."